Required:

  1. Firm A has a margin of 11%, sales of $560,000, and ROI of 18%. Calculate the firm's average total assets. (Round "Turnover" to 1 decimal place.)
  2. Firm B has net income of $74,000, turnover of 1.20, and average total assets of $900,000. Calculate the firm's sales, margin, and ROI. (Round "Margin" and "ROI" answers to 1 decimal place.)
  3. Firm C has net income of $134,000, turnover of 2.01, and ROI of 23.40%. Calculate the firm's margin, sales, and average total assets. (Round "Margin" answer to 1 decimal place and use the rounded answer for the subsequent computations.)
Answer #1

Answers

--With working and calculation

Firm A
A Sales $560,000
B Margin 11%
C = A x B Net Income $61,600
D ROI 18%
E = D/B Turnover ratio 1.6
F = A/E Average Total Assets $350,000 Answer
Firm B
A Net Income $74,000
B Turnover 1.2
C Average Total Assets $900,000
D = B x C Sales $1,080,000 Answer
E = A/D Margin 6.9% Answer
F = B x E ROI 8.3% Answer
Firm C
A Net Income $134,000
B Turnover 2.01
C ROI 23.4%
D = C/B Margin 11.6% Answer
F = A/D Sales $1,155,172 Answer
G = F/B Average Total Assets $574,712 Answer